# Exploring Kinetic Curves Features for the Classification of Benign and Malignant Breast Lesions in DCE-MRI

### v1

@@ Abstract
Breast cancer is the most common malignant tumor among women and the second cause of cancer-related death. Early diagnosis in clinical practice is crucial for timely treatment and prognosis. Dynamic contrast-enhanced magnetic resonance imaging (DCE-MRI) has revealed great usability in the preoperative diagnosis and assessing therapy effects thanks to its capability to reflect the morphology and dynamic characteristics of breast lesions. However, most existing computer-assisted diagnosis algorithms only consider conventional radiomic features when classifying benign and malignant lesions in DCE-MRI. In this study, we propose to fully leverage the dynamic characteristics from the kinetic curves as well as the radiomic features to boost the classification accuracy of benign and malignant breast lesions. The proposed method is a fully automated solution by directly analyzing the 3D features from the DCE-MRI. The proposed method is evaluated on an in-house dataset including 200 DCE-MRI scans with 298 breast tumors (172 benign and 126 malignant tumors), achieving favorable classification accuracy with an area under curve (AUC) of 0.94. By simultaneously considering the dynamic and radiomic features, it is beneficial to effectively distinguish between benign and malignant breast lesions.
@@ II-A Automated Tumor Segmentation
To conduct the 3D tumor segmentation from DCE-MRI scans, we employ our previously developed joint-phase attention (JPA) network [16]. The JPA network is trained using a large-scale multi-center breast DCE-MRI dataset thus is with favorable segmentation accuracy and generalization performance. By using the JPA network, we can obtain the segmentation results of the whole breast and the tumor region.
@@ III-A Materials
The proposed method was evaluated using an in-house breast DCE-MRI dataset collected from the Cancer Center of Sun Yat-Sen University. The whole dataset included 200 DCE-MRI scans with 298 biopsy-proven breast tumors (172 benign tumors and 126 malignant tumors). Five-fold cross validation was conducted to investigate the classification performance.
@@ III-C Feature Selection Results
Features Used Accuracy Recall Precision F1 score AUC
Dynamic Features 0.7471 0.7583 0.7109 0.7339 0.8253
Radiomic Features 0.8352 0.8917 0.7810 0.8327 0.9080
Combined Features 0.8889 0.9417 0.8370 0.8863 0.9476
@@ IV Conclusion
In this study, we explore the practicality of dynamic features and radiomic features to distinguish between benign and malignant breast lesions in DCE-MRI. To this end, we design a fully automated solution to analyze the 3D dynamic and radiomic features, and use them to construct the classification model. The experimental results on an in-house DCE-MRI dataset show the efficacy of the proposed method. By simultaneously considering the dynamic and radiomic features, it is beneficial to effectively distinguish between benign and malignant breast lesions.

### v2

@@ Abstract
Breast cancer is the most common malignant tumor among women and the second cause of cancer-related death. Early diagnosis in clinical practice is crucial for timely treatment and prognosis. Dynamic contrast-enhanced magnetic resonance imaging (DCE-MRI) has revealed great usability in the preoperative diagnosis and assessing therapy effects thanks to its capability to reflect the morphology and dynamic characteristics of breast lesions. However, most existing computer-assisted diagnosis algorithms only consider conventional radiomic features when classifying benign and malignant lesions in DCE-MRI. In this study, we propose to fully leverage the dynamic characteristics from the kinetic curves as well as the radiomic features to boost the classification accuracy of benign and malignant breast lesions. The proposed method is a fully automated solution by directly analyzing the 3D features from the DCE-MRI. The proposed method is evaluated on an in-house dataset including 200 DCE-MRI scans with 298 breast tumors (172 benign and 126 malignant tumors), achieving favorable classification accuracy with an area under curve (AUC) of 0.94. By simultaneously considering the dynamic and radiomic features, it is beneficial to effectively distinguish between benign and malignant breast lesions. The algorithm is publicly available at https://github.com/ryandok/JPA.
@@ II-A Automated Tumor Segmentation
To conduct the 3D tumor segmentation from DCE-MRI scans, we employ our previously developed joint-phase attention (JPA) network [16]. The JPA network is trained using a large-scale multi-center breast DCE-MRI dataset thus is with favorable segmentation accuracy and generalization performance. By using the JPA network, we can obtain the segmentation results of the whole breast and the tumor region (denoted as $tumor$ $mask$).
Figure 2: Different types of the kinetic curves.
@@ III-A Materials
This retrospective study was conducted with the approval of the Institutional Review Board of the Cancer Center of Sun Yat-Sen University, and the necessity for obtaining written informed consent was waived. The whole collected dataset included 200 DCE-MRI scans (patients) with 298 biopsy-proven breast tumors (172 benign tumors and 126 malignant tumors). Five-fold cross validation was conducted to investigate the classification performance.
@@ III-B Evaluation Metrics
Features Used Accuracy Recall Precision F1 score AUC
Dynamic Features 0.7471 0.7583 0.7109 0.7339 0.8253
Radiomic Features 0.8352 0.8917 0.7810 0.8327 0.9080
Combined Features 0.8889 0.9417 0.8370 0.8863 0.9476
@@ IV Conclusion
In this study, we explore the practicality of dynamic features and radiomic features to distinguish between benign and malignant breast lesions in DCE-MRI. To this end, we design a fully automated solution to analyze the 3D dynamic and radiomic features, and use them to construct the classification model. The experimental results on an in-house DCE-MRI dataset show the efficacy of the proposed method. By simultaneously considering the dynamic and radiomic features, it is beneficial to effectively distinguish between benign and malignant breast lesions. It is worth noting that the performance and practicability of the computer assisted diagnosis algorithms can still be enhanced. On one hand, the 3D tumor segmentation methods can be more accurate and annotation-efficient, by further employing recently developed attention mechanisms [25, 26], or weakly-/un-supervised segmentation methods [27, 28, 29]. On the other hand, the classification method used in this study is the simple LDA model, while future studies may consider leverage other classification algorithms [30]. In addition, deep learning methods that directly analyze DCE-MRI scans and output diagnosis results can be investigated.